OpenStudy (istim):

Determine dy/dx for y=5sqrt(x^3)

OpenStudy (anonymous):

y=5sqrt(x^3) = 5(x^(3/2)) dy/dx = 5(3/2)(x^(3/2 -1)) = 15/2 (x^1/2)

OpenStudy (anonymous):

IsTim now you're here for math too? giggle, shiny and math?

OpenStudy (anonymous):

y=x^n dy/dx = nx^(n-1) this is the property I use

OpenStudy (istim):

Truly, I don't really understand what derivatives is. @Ishaan94 Pardon?

OpenStudy (istim):

The Power Rule. Ok.

OpenStudy (anonymous):

Sorry.

OpenStudy (istim):

It's ok. I don't understand (your speech and derivatives; either/or) though.

OpenStudy (istim):

If you can, please describe each step thoroughly. Thank you.

OpenStudy (istim):

OH well.

OpenStudy (istim):

What's happening at dy/dx = 5(3/2)(x^(3/2 -1)) ?

OpenStudy (istim):

Or, how did it come to that?

OpenStudy (anonymous):

y=ax^n dy/dx =a( nx^(n-1)) put a=1 and n=3/2

OpenStudy (istim):

Uhm, sorry to ask, but could you type dy/dx = 5(3/2)(x^(3/2 -1)) in Latex (The Equation Editor)? I don't think I'm understanding it correctly.

OpenStudy (istim):

@CoCoTsoi When you have time, of course.

OpenStudy (anonymous):